OpenStudy (anonymous):

Prove (1+Cosx)/sinx = sinx /( 1-cosx) 1st what do i need to take note??

OpenStudy (anonymous):

Do you want to add any brackets? So far, it says this: \[1 + \frac{\cos x}{\sin x} = \frac{\sin x}{1} - \cos x\]

OpenStudy (anonymous):

Thanks for editing the post and adding the brackets...

OpenStudy (anonymous):

note: multiply top and bottom by \(1-\cos x\) and note the pythagorean identity.

OpenStudy (anonymous):

First thing: multiply both sides by certain values to remove the fractions. Eg: \[\frac{a}{b} = \frac{c}{d} \implies ad = bc\] Do that.

OpenStudy (raden):

for left side, if multiply it by conyugate of numerator gives (1+Cosx)/sinx * (1-cosx)/(1-cosx) = (1-cos^2 x)/(sinx*(1-cosx)) = sin^2 x / (sinx*(1-cosx)) simplify again...

OpenStudy (anonymous):

@scarydoor you cannot do that. this is a RHS--> LHS proof.

OpenStudy (anonymous):

@Shadowys yes, you can do that.

OpenStudy (anonymous):

............ you can't use something you haven't proved yet.

OpenStudy (nubeer):

i dont think we can cross multiply while prooving @scarydoor

OpenStudy (anonymous):

By using equivalently true statements, you can. You just need to be careful that the order of implications runs the right way. Above, I should have said "iff" not "implies", to avoid your complaint.

OpenStudy (anonymous):

\[\frac{a}{b}=\frac{c}{d} \iff ad = bc.\]

OpenStudy (anonymous):

you're trying to prove \(\frac{1+\cos x}{\sin x} = \frac{\sin x}{1-\cos x} \) you don't know if \(\frac{1+\cos x}{\sin x} = \frac{\sin x}{1-\cos x} \) if and only if\(\frac{1+\cos x}{\sin x} = \frac{\sin x}{1-\cos x} \) then \(\frac{1+\cos x}{\sin x} = \frac{\sin x}{1-\cos x} \)

OpenStudy (anonymous):